OpenStudy (anonymous):

cost of truck is $18,000 and approved for 8% loan.What would the additional cost be if you choose a 60 month term instead of 48 month, which one would be better? Loan Amortization table at 8% is for 48 months: 24.42 and for 60 months 20.28 Can anyone help?

OpenStudy (tkhunny):

You have the payments and the terms. What's stopping you from finishing the calculations? 48*24.42 = ?? 60*20.28 = ??

OpenStudy (anonymous):

I was not sure of the formula...

OpenStudy (tkhunny):

What formula? The amortization payment is already calculated and given in the problem statement. 48*24.42 = ?? 60*20.28 = ??

OpenStudy (anonymous):

but how does that fit in with the rest of the problem, 18,000 at 8%

OpenStudy (tkhunny):

You are looking beyond the mark. 18000 is of little additional consequence. The interest rate will not help you answer the question. You already used the interest rate to look up the payments. Total Paid on the 48 month term 48*24.42 = ?? Total paid on the 60 month term 60*20.28 = ?? Calculate these two values and then answer the question. You will have to understand your table. Is that the monthly payment per $1000 loaned? If so, multiply your results by 18 and you are done.

OpenStudy (anonymous):

ok, so I have for additional cost of 48 months= 1,172.16 60months= 1,216.80

OpenStudy (tkhunny):

That's not the "additional" cost. That's the "Cost per $1000 borrowed". Now, calculate the full amounts and subtract.